OpenStudy (anonymous):

the table list data regarding the average salaries of professional athletes in the years 1991 and 2001 1991 262,000 2001 1,390,000 a: use the data points to find a linear function that fits the data b:use the function to predict the average salary in 2005 and 2010 S(x)=?

OpenStudy (anonymous):

if we set 1991 as year 0 and set 2001 as year 10 we get S(0) = 262 000 S(10) = 1 390 000 for a linear function we have S(x) = a * x + b for the first it is easy to see S(0) = a * 0 + b = 262 000 b = 262 000 and the second gives S(10) = a * 10 + 262 000 = 1 390 000 a * 10 = 1 390 000 - 262 000 => a * 10 = 1 128 000 a = 112 800 S(x) = 112 800 * x + 262 000 then insert for x = 14 for year 2005 S(14) = 112 800 * 14 + 262 000 S(14) = 1 841 200 and x = 19 for year 2010 S(19) = 112 800 * 19 + 262 000 S(19) = 2 405 200